Randy and his sister agreed that they would each mow one-half of the lawn. The lawn is a 25' by 45' rectangle. The lawn mower cuts a path two feet wide. If Randy starts at one corner and mows a path completely around the outside, how many times should he go around to mow one-half of the lawn?
Answer by ankor@dixie-net.com(22740)

Randy and his sister agreed that they would each mow one-half of the lawn. The lawn is a 25' by 45' rectangle. The lawn mower cuts a path two feet wide. If Randy starts at one corner and mows a path completely around the outside, how many times should he go around to mow one-half of the lawn?
:
Find the area of the lawn: 25 * 45 = 1125 sq/ft
Half the area = 562.5 sq/ft
:
Find the perimeter: 2(25) + 2(45) = 140 ft
The first round he will mow: 140 * 2 = 280 sq/ft
:
He has subtracted 4 ft from each dimension now the perimeter is:
2(21) + 2(41) = 124
2nd round he will mow: 124 * 2 = 248 sq/ft
:
Total area mowed now: 280 + 248 = 528 sq/ft
:
That leaves 562.5 - 528 = 34.5 sq/ft to mow
:
He has subtracted another 4 ft from each dimension, now the perimeter is:
2(17) + 2(37) = 108 ft
3rd round he will mow 108 * 2 = 216 sq/ft
:
But he only needs to mow 34.5 sq/ft. If mows about 17 ft (1 side) of the 3rd round he will have it.
:
We can say 2 rounds + plus 1 side will be about half